# Near-Optimal Regret for Efficient Stochastic Combinatorial Semi-Bandits

### v1

@@ Abstract
The combinatorial multi-armed bandit (CMAB) is a cornerstone of sequential decision-making framework, dominated by two algorithmic families: UCB-based and adversarial methods such as follow the regularized leader (FTRL) and online mirror descent (OMD). However, prominent UCB-based approaches like CUCB suffer from additional regret factor $\log T$ that is detrimental over long horizons, while adversarial methods such as EXP3.M and HYBRID impose significant computational overhead. To resolve this trade-off, we introduce the Combinatorial Minimax Optimal Strategy in the Stochastic setting (CMOSS). CMOSS is a computationally efficient algorithm that achieves an instance-independent regret of $O\big{(}(\log k)^{2}\sqrt{kmT}\big{)}$ under semi-bandit feedback, where $m$ is the number of arms and $k$ is the maximum cardinality of a feasible action. Crucially, this result eliminates the dependency on $\log T$ and matches the established $\Omega\big{(}\sqrt{kmT}\big{)}$ lower bound up to $O\big{(}(\log k)^{2}\big{)}$. We then extend our analysis to show that CMOSS is also applicable to cascading feedback. Experiments on synthetic and real-world datasets validate that CMOSS consistently outperforms benchmark algorithms in both regret and runtime efficiency.
@@ 1 Introduction
The stochastic multi-armed bandit (MAB) is a fundamental $T$-round sequential decision-making game that has been widely studied in the past two decades (Audibert and Bubeck 2009; Auer et al. 1995, 2002; Robbins 1952). In this game, one player faces $m$ arms. In each round $t\in[T]$, the player chooses an arm $A_{t}$ among the $m$ arms and gets a reward drawn from a predetermined but unknown reward distribution with mean $\mu_{A_{t}}\in[0,1]$. The arm with the largest reward mean $\mu_{*}$ is called the best arm. The total expected regret is defined as $\mathbf{E}\left[R(T)\right]=\mathbf{E}\left[\sum_{t=1}^{T}\left(\mu_{*}-\mu_{A_{t}}\right)\right],$ where the expectation is over the randomness from the player’s strategy and the random reward outcome. And the aim is to minimize the minimax regret, namely the regret of the best algorithm against the worst input.
@@
In many real-world applications, however, decision-making problems often extend beyond the simple MAB framework to involve combinatorial actions among multiple arms. For instance, in online advertising (Kveton et al. 2015a), an advertiser could select up to $k$ web pages to display ads, constrained by a budget. Each user visits a subset of pages, and the advertiser must learn the unknown click-through probabilities to maximize clicks. Similarly, in viral marketing (Gai, Krishnamachari, and Jain 2012; Tang et al. 2017), a marketer selects seed nodes in a social network to initiate information cascades, learning influence probabilities to maximize the reach of campaigns. These scenarios highlight the need for algorithms that efficiently address the combiantorial MAB (CMAB) problem.
While a number of works (Combes et al. 2015; Kveton et al. 2015c; Chen et al. 2016; Wang and Chen 2017; Liu et al. 2022, 2024, 2025) have focused on the CMAB problem, existing algorithms either suffer from additional $\log T$ terms in their regret bounds or incur high computational costs in practice. Among them, the Combinatorial UCB (CUCB) algorithm (Chen, Wang, and Yuan 2013; Kveton et al. 2015c; Wang and Chen 2017; Liu et al. 2022) is one of the most prominent and widely adopted approaches. The CUCB algorithm extends the classic UCB framework by selecting the action (super arm) from a feasible set that solves a combinatorial optimization problem defined over the upper confidence bounds of the base arms. Kveton et al. (Kveton et al. 2015c) establish an upper regret bound of $O\left(\sqrt{kmT\log T}\right)$ for CUCB, which matches the lower bound of $\Omega\left(\sqrt{kmT}\right)$ only up to a $\sqrt{\log T}$ factor.
@@ 1.1 Our results
To address the above limitation, we study the stochastic combinatorial semi-bandit problem and propose the CMOSS algorithm. Our contribution is to show that CMOSS is not only computationally efficient but also has a regret of $O\left((\log k)^{2}\sqrt{kmT}\right)$. We also extend the setting to cascading feedback and compare our result with CUCB, since the other three algorithms can not be applied to the scenarios where arms in $A_{t}$ are probabilistically triggered. The results are summarized in Table 1.
[…]
Computation Complexity
Per-Round (Average,
$10^{-5}$ sec) $\dagger$
COMBEXP $O\left(\sqrt{k^{3}mT\log m}\right)$ None Too large
CUCB $O\left(\sqrt{kmT\log T}\right)$ $O\left(\sqrt{kmT\log T}\right)$ 5.743
EXP3.M $O\left(\sqrt{kmT\log\left(\frac{m}{k}\right)}\right)$ None 43.095
[…]
$O\left((m-k)\sqrt{\log\left(\frac{m}{m-k}\right)T}\right)$, $k>\frac{m}{2}$ None None
CMOSS $O\left((\log k)^{2}\sqrt{kmT}\right)$ $O\left(\frac{(\log k)^{2}}{p^{*}}\sqrt{kmT}\right)$ 6.706
Table 1: Theoretical regret bounds and empirical computational cost per-round for algorithms on the CMAB problem.
@@
$\dagger$ We fix $k=10$, $m=30$, and $T=100000$ for runtime measurement on a synthetic dataset (averaged over 10 runs) in the semi-bandit feedback setting. For HYBRID, we capture the intuition of the computation complexity through the case when $k\leq\frac{m}{2}$. See Section 4.1 for dataset details.
$\ddagger$ Although when $k\leq\frac{m}{2}$ the regret bound is optimal, HYBRID performs relatively bad in experiments simulating realistic conditions where arm means are in the range of $[0,0.1]$. See Section 4.1 for dataset details.
@@
Furthermore, we show that CMOSS could also be extended to cascading feedback. More details are given in Section 3.3. We also run comparative experiments evaluating the performance of CUCB and CMOSS in Appendix C.4.
We also conduct comparative experiments involving CMOSS, CUCB, EXP3.M, and HYBRID on both synthetic and real-world datasets under semi-bandit feedback, including ablation studies. Detailed descriptions of CUCB, EXP3.M, and HYBRID are provided in Appendix B for completeness. The results demonstrate that, under noisy conditions and after sufficient exploration rounds, CMOSS incurs slightly higher computational complexity than CUCB while achieving notably lower cumulative regret. Compared to EXP3.M and HYBRID, CMOSS attains better cumulative regret with significantly reduced computational complexity. This performance profile highlights CMOSS’s superior balance between exploration and exploitation, underscoring its effectiveness in such environments. Further details are provided in Section 4.
@@ 1.2 Related work
Stochastic Combinatorial Multi-Armed Bandits (CMAB). The stochastic CMAB problem has been extensively studied in the literature (Chen et al. 2016; Chen, Wang, and Yuan 2013; Combes et al. 2015; Kveton et al. 2015a, c; Wang and Chen 2017; Liu et al. 2022, 2023). The seminal work by Gai et al. (Gai, Krishnamachari, and Jain 2012) pioneered the study of stochastic CMAB with semi-bandit feedback, which inspired algorithmic advancements and improved regret bound across various settings (Kveton et al. 2015a, 2014; Tang et al. 2017). For the instance-dependent regret, Wang et al. (Wang and Chen 2018) proposed the combinatorial Thompson sampling (CTS) algorithm with an offline oracle, achieving a regret of $O\left(\frac{m\log T}{\Delta_{\min}}\right)$, where $\Delta_{min}$ is the minimum gap between the expected reward of the optimal action and any suboptimal action. The bound essentially matches asymptotically on $T$.
For the instance-independent regret, one line of works focused on the adversarial setting and designed algorithms based on the principle of OMD and FTRL. Several studies (Cesa-Bianchi and Lugosi 2012; Dani, Kakade, and Hayes 2007; Uchiya, Nakamura, and Kudo 2010; Zimmert, Luo, and Wei 2019) proposed enhancements such as COMBEXP, EXP3.M, with the best (Zimmert, Luo, and Wei 2019) achieving a regret of $O\left(\sqrt{kmT}\right)$ when $k\leq\frac{m}{2}$ and $O\left((m-k)\sqrt{\log\left(\frac{m}{m-k}\right)T}\right)$ when $k>\frac{m}{2}$. However, these methods suffer from high computational overhead. Another line of works (Chen, Wang, and Yuan 2013; Liu et al. 2022; Wang and Chen 2017, 2018; Liu et al. 2023) used UCB-based methods. Chen et al. (Chen, Wang, and Yuan 2013) proposed the CUCB algorithm and analyzed the regret of CUCB with an approximation oracle. Then Kveton et al. (Kveton et al. 2015c) derived an upper bound of $O\left(\sqrt{kmT\log T}\right)$ for CUCB and established a lower bound of $\Omega\left(\sqrt{kmT}\right)$. To address the degradation caused by the $\log T$ term over long horizons, we refine the upper confidence bound and propose the CMOSS algorithm. It achieves strong performance while maintaining low computational costs, even after removing the $\log T$ term.
@@ 2 Preliminaries
We begin by introducing the general framework of the stochastic CMAB and establishing notations used throughout the paper. The CMAB problem is modeled as a sequential decision-making game between a learning agent and an environment. The environment consists of $m$ independent base arms $\mathtt{arm}_{1},\ldots,\mathtt{arm}_{m}$, each associated with a random reward $X_{i}\sim D_{i}$ supported on $[0,1]$ and having mean $\mu_{i}$. The distributions $\{D_{i}\}_{i=1}^{m}$ are fixed by the environment before the game begins and are drawn from a known class $\mathcal{D}$. The learner knows $\mathcal{D}$ but not the specific instantiation of each $D_{i}$.
The action space $\mathcal{A}$ consists of all subsets of $[m]$ such that $\forall A\in\mathcal{A}$, $\left|A\right|\leq k$. The $T$-round decision game starts as follows: in each round $t\in[T]$, the player selects an action $A_{t}$ from $\mathcal{A}$ based on the feedback history, and the environment draws from every distribution $D_{i}$ a sample $X^{(t)}_{i}$. Let $\bm{\mu}$ represents the vector of the mean rewards of all arms. After the action $A_{t}$ is played, under semi-bandit feedback, the outcomes of $X^{(t)}_{i}$ for all $i\in A_{t}$ are observed as the feedback to the player. Under cascading feedback, the arms in $A_{t}$ are probabilistically triggered and only the arms in the triggered set $\tau_{t}$ could be observed. In both cases the player receives $r(A_{t},\bm{\mu})$ as the reward in round $t$.
@@ 3.1 The CMOSS algorithm
The CMOSS algorithm maintains the empirical estimate $\hat{\mu}_{i}$ for the true mean $\mu_{i}$ for all the $m$ arms, then selects the action by solving the combinatorial problem of maximizing $r(A,\bar{}\bm{\mu})$, where each $\bar{\mu}_{i}$ optimistically estimates $\mu_{i}$ by a confidence interval $\rho_{i}$. The prominent difference between CMOSS and CUCB is that the radius $\rho_{i}=\sqrt{\frac{1}{T_{i}}\ln^{+}\left(\frac{1}{\delta T_{i}}\right)}$ but not $\sqrt{\frac{3\ln t}{2T_{i}}}$, where $\delta$ is a parameter that scales in proportion to $\frac{1}{T}$. In this case, the probability that $\hat{\mu}_{i}$ lies in the confidence region $(\mu_{i}-\rho_{i},\mu_{i}+\rho_{i})$ in every round $t$ could be much less than $\Theta\left(\frac{1}{t^{2}}\right)$. Thus we can not neglect the regret caused by $\hat{\mu}_{i}$ falling outside $(\mu_{i}-\rho_{i},\mu_{i}+\rho_{i})$. Therefore we need to bound the regret in the average sense. We have the following theorem:
@@
For any stochastic combinatorial semi-bandit feedback instance with $m$ arms, time horizon $T$, and the maximum number of arms that could be chosen in a round is $k$, then the expected regret of Algorithm 1 is $O\left((\log k)^{2}\sqrt{kmT}\right)$.
Theorem 1 eliminates the $\sqrt{\log T}$ factor in the upper bound of CUCB. Unlike the way EXP3.M and HYBRID do, CMOSS also retains the low computation complexity, introducing only an additional $O\left(\left(\log k\right)^{2}\right)$ term in the regret bound.

### v2

@@ Abstract
The combinatorial multi-armed bandit (CMAB) is a cornerstone of sequential decision-making framework, dominated by two algorithmic families: UCB-based and adversarial methods such as follow the regularized leader (FTRL) and online mirror descent (OMD). However, prominent UCB-based approaches like CUCB suffer from additional regret factor $\log T$ that is detrimental over long horizons, while adversarial methods such as EXP3.M and HYBRID impose significant computational overhead. To resolve this trade-off, we introduce the Combinatorial Minimax Optimal Strategy in the Stochastic setting (CMOSS). CMOSS is a computationally efficient algorithm that achieves an instance-independent regret of $O\big((\log k)\sqrt{kmT}\big)$ when $k\leq\frac{m}{2}$ and $O\big((m-k)\sqrt{\log k\log(m-k)T}\big)$ when $k>\frac{m}{2}$ under semi-bandit feedback, where $m$ is the number of arms and $k$ is the maximum cardinality of a feasible action. Crucially, this result eliminates the dependency on $\log T$ and matches the established lower bounds of $\Omega\big(\sqrt{kmT}\big)$ when $k\leq\frac{m}{2}$ and $\Omega\big((m-k)\sqrt{\log(\frac{m}{m-k})T}\big)$ when $k>\frac{m}{2}$ up to logarithmic terms of $k$ and $m$. We then extend our analysis to show that CMOSS is also applicable to cascading feedback. Experiments on synthetic and real-world datasets validate that CMOSS consistently outperforms benchmark algorithms in both regret and runtime efficiency.
@@ 1 Introduction
The stochastic multi-armed bandit (MAB) is a fundamental $T$-round sequential decision-making game that has been widely studied in the past two decades (Audibert and Bubeck 2009; Auer et al. 1995, 2002; Robbins 1952). In this game, one player faces $m$ arms. In each round $t\in[T]$, the player chooses an arm $A_{t}$ among the $m$ arms and gets a reward drawn from a predetermined but unknown reward distribution with mean $\mu_{A_{t}}\in[0,1]$. The arm with the largest reward mean $\mu_{*}$ is called the best arm. The total expected regret is defined as $\mathbf{E}\left[R(T)\right]=\mathbf{E}\left[\sum_{t=1}^{T}\left(\mu_{*}-\mu_{A_{t}}\right)\right],$ where the expectation is over the randomness from the player’s strategy and the random reward. And the objective is to minimize the minimax regret, namely the regret of the best algorithm against the worst input.
@@
In many real-world applications, however, decision-making problems often extend beyond the simple MAB framework to involve combinatorial actions among multiple arms. For instance, in online advertising (Kveton et al. 2015a), an advertiser could select up to $k$ web pages to display ads, constrained by a budget. Each user visits a subset of pages, and the advertiser must learn the unknown click-through probabilities to maximize clicks. Similarly, in viral marketing (Gai, Krishnamachari, and Jain 2012; Tang et al. 2017), a marketer selects seed nodes in a social network to initiate information cascades, learning influence probabilities to maximize the reach of campaigns. These scenarios highlight the need for algorithms that efficiently address the combinatorial MAB (CMAB) problem.
While a number of works (Combes et al. 2015; Kveton et al. 2015c; Chen et al. 2016; Wang and Chen 2017; Liu et al. 2022, 2024, 2025) have focused on the CMAB problem, existing algorithms either suffer from additional $\log T$ terms in their regret bounds or incur high computational costs in practice. Among them, the Combinatorial UCB (CUCB) algorithm (Chen, Wang, and Yuan 2013; Kveton et al. 2015c; Wang and Chen 2017; Liu et al. 2022) is one of the most prominent and widely adopted approaches. The CUCB algorithm extends the classic UCB framework by selecting the action (super arm) from a feasible set that solves a combinatorial optimization problem defined over the upper confidence bounds of the base arms. Kveton et al. (Kveton et al. 2015c) establish an upper regret bound of $O\left(\sqrt{kmT\log T}\right)$ for CUCB, which matches the lower bound of $\Omega\left(\sqrt{kmT}\right)$ only up to a $\sqrt{\log T}$ factor when $k\leq\frac{m}{2}$.
@@ 1.1 Our results
To address the above limitation, we study the stochastic combinatorial semi-bandit problem and propose the CMOSS algorithm. Our contribution is to show that CMOSS is not only computationally efficient but also has a regret of $O\big((\log k)\sqrt{kmT}\big)$ when $k\leq\frac{m}{2}$ and $O\big((m-k)\sqrt{\log k\log(m-k)T}\big)$ when $k>\frac{m}{2}$. We also extend the setting to cascading feedback and compare our result with CUCB, since the other three algorithms can not be applied to the scenarios where arms in $A_{t}$ are probabilistically triggered. The results are summarized in Table 1.
Table 1: Regret bounds and empirical computational cost per-round for algorithms on the CMAB problem.
@@
$\dagger$ We select $k=20,m=30$ for HYBRID and CMOSS when $k>\frac{m}{2}$ and $k=10,m=30$ for other rows with $T=100000$ separately for runtime measurement on a synthetic dataset (averaged over 10 runs) in the semi-bandit feedback setting. See Section 4.1 for dataset details.
$\ddagger$ HYBRID performs relatively bad in experiments simulating realistic conditions where arm means are in the range of $[0,0.1]$, although the regret bound is optimal. See Section 4.1 for dataset details.
[…]
COMBEXP $O\left(\sqrt{k^{3}mT\log m}\right)$ None Too large
CUCB $O\left(\sqrt{kmT\log T}\right)$ $O\left(\sqrt{kmT\log T}\right)$ 1.516
EXP3.M $O\left(\sqrt{kmT\log\left(\frac{m}{k}\right)}\right)$ None 10.429
[…]
$O\left((m-k)\sqrt{\log\left(\frac{m}{m-k}\right)T}\right)$, $k>\frac{m}{2}$ None 11.195
CMOSS $O\left((\log k)\sqrt{kmT}\right)$, $k\leq\frac{m}{2}$ $O\left(\frac{\log k}{p^{*}}\sqrt{kmT}\right)$ 1.737
$O\left((m-k)\sqrt{\log k\log(m-k)T}\right)$, $k>\frac{m}{2}$ $O\left(\frac{m-k}{p^{*}}\sqrt{\log k\log(m-k)T}\right)$ 3.067
@@
Furthermore, we show that CMOSS could also be extended to cascading feedback. More details are given in Section 3.3. We also run comparative experiments evaluating the performance of CUCB and CMOSS in Section C.4.
We also conduct comparative experiments involving CMOSS, CUCB, EXP3.M, and HYBRID on both synthetic and real-world datasets under semi-bandit feedback, including ablation studies. Detailed descriptions of CUCB, EXP3.M, and HYBRID are provided in Appendix B for completeness. The results demonstrate that, under noisy conditions and after sufficient exploration rounds, CMOSS incurs slightly higher computational cost than CUCB while achieving notably lower cumulative regret. Compared to EXP3.M and HYBRID, CMOSS attains better cumulative regret with significantly reduced computational cost. This performance profile highlights CMOSS’s superior balance between exploration and exploitation, underscoring its effectiveness in such environments. Further details are provided in Section 4.
@@ 1.2 Related work
Stochastic Combinatorial Multi-Armed Bandits (CMAB). The stochastic CMAB problem has been extensively studied in the literature (Chen et al. 2016; Chen, Wang, and Yuan 2013; Combes et al. 2015; Kveton et al. 2015a, c; Wang and Chen 2017; Liu et al. 2022, 2023). The seminal work by Gai et al. (Gai, Krishnamachari, and Jain 2012) pioneered the study of stochastic CMAB with semi-bandit feedback, which inspired algorithmic advancements and improved regret bound across various settings (Kveton et al. 2015a, 2014; Tang et al. 2017). For the instance-dependent regret, Wang et al. (Wang and Chen 2018) proposed the combinatorial Thompson sampling (CTS) algorithm with an offline oracle, achieving a regret of $O\left(\frac{m\log T}{\Delta_{\min}}\right)$, where $\Delta_{min}$ is the minimum gap between the expected reward of the optimal action and any suboptimal action. The bound essentially matches asymptotically on $T$. Thus in this paper, we focus on the instance-independent regret.
For the instance-independent regret, one line of works focused on the adversarial setting and designed algorithms based on the principle of OMD and FTRL. Several studies (Cesa-Bianchi and Lugosi 2012; Dani, Kakade, and Hayes 2007; Uchiya, Nakamura, and Kudo 2010; Zimmert, Luo, and Wei 2019) proposed enhancements such as COMBEXP, EXP3.M, with the best (Zimmert, Luo, and Wei 2019) achieving a regret of $O\left(\sqrt{kmT}\right)$ when $k\leq\frac{m}{2}$ and $O\left((m-k)\sqrt{\log\left(\frac{m}{m-k}\right)T}\right)$ when $k>\frac{m}{2}$. The latter case matches the lower bound of $\Omega\left((m-k)\sqrt{\log\left(\frac{m}{m-k}\right)T}\right)$. However, these methods suffer from high computational overhead. Another line of works (Chen, Wang, and Yuan 2013; Liu et al. 2022; Wang and Chen 2017, 2018; Liu et al. 2023) used UCB-based methods. Chen et al. (Chen, Wang, and Yuan 2013) proposed the CUCB algorithm and analyzed the regret of CUCB with an approximation oracle. Then Kveton et al. (Kveton et al. 2015c) derived an upper bound of $O\left(\sqrt{kmT\log T}\right)$ for CUCB and established a lower bound of $\Omega\left(\sqrt{kmT}\right)$ when $k\leq\frac{m}{2}$. To address the degradation caused by the $\log T$ term over long horizons, we refine the upper confidence bound and propose the CMOSS algorithm. It achieves strong performance while maintaining low computational costs, even after removing the $\log T$ term.
@@ 2 Preliminaries
We begin by introducing the general framework of the stochastic CMAB and establishing notations used throughout the paper. The CMAB problem is modeled as a sequential decision-making game between a learning agent and an environment. The environment consists of $m$ base arms $\mathtt{arm}_{1},\ldots,\mathtt{arm}_{m}$, each associated with a random reward $X_{i}\sim D_{i}$ supported on $[0,1]$ and having mean $\mu_{i}$. The distributions $\{D_{i}\}_{i=1}^{m}$ are fixed by the environment before the game begins and are drawn from a known class $\mathcal{D}$. The learner knows $\mathcal{D}$ but not the specific instantiation of each $D_{i}$. In each round, the rewards of these arms may be correlated with each other.
The action space $\mathcal{A}$ consists of all subsets of $[m]$ such that $\forall A\in\mathcal{A}$, $\left|A\right|\leq k$. The $T$-round decision game starts as follows: in each round $t\in[T]$, the player selects an action $A_{t}$ from $\mathcal{A}$ based on the feedback history, and the environment draws from every distribution $D_{i}$ a sample $X^{(t)}_{i}$. Let $\bm{\mu}$ represents the vector of the mean rewards of all arms. After the action $A_{t}$ is played, under semi-bandit feedback, the values of $X^{(t)}_{i}$ for all $i\in A_{t}$ are observed as the feedback to the player. Under cascading feedback, the arms in $A_{t}$ are probabilistically triggered and only the arms in the triggered set $\tau_{t}$ could be observed. In both cases the player receives $r(A_{t},\bm{\mu})$ as the reward in round $t$.
@@ 3.1 The CMOSS algorithm
The CMOSS algorithm maintains the empirical estimate $\hat{\mu}_{i}$ for the true mean $\mu_{i}$ for all the $m$ arms, then selects the action by solving the combinatorial problem of maximizing $r(A,\bar{}\bm{\mu})$, where each $\bar{\mu}_{i}$ optimistically estimates $\mu_{i}$ by a confidence interval $\rho_{i}$. The prominent difference between CMOSS and CUCB is that the radius $\rho_{i}=\sqrt{\frac{1}{T_{i}}\ln^{+}\left(\frac{1}{\delta T_{i}}\right)}$ but not $\sqrt{\frac{3\ln t}{2T_{i}}}$, where $\delta$ is a parameter to be optimized during the proof. In this case, the probability that $\hat{\mu}_{i}$ lies in the confidence region $(\mu_{i}-\rho_{i},\mu_{i}+\rho_{i})$ in every round $t$ could be much less than $\Theta\left(\frac{1}{t^{2}}\right)$. Thus we can not neglect the regret caused by $\hat{\mu}_{i}$ falling outside $(\mu_{i}-\rho_{i},\mu_{i}+\rho_{i})$. Therefore we need to bound the regret in the average sense. We have the following theorem:
@@
For any stochastic combinatorial semi-bandit feedback instance with $m$ arms, time horizon $T$, and the maximum number of arms that could be chosen in a round is $k$, then when $k\leq\frac{m}{2}$, the expected regret of Algorithm 1 is $O\left((\log k)\sqrt{kmT}\right)$ and when $k>\frac{m}{2}$, the expected regret is $O(\left(m-k\right)\sqrt{\log k\log(m-k)T})$.
Theorem 1 eliminates the $\sqrt{\log T}$ factor in the upper bound of CUCB. Unlike the way EXP3.M and HYBRID do, CMOSS also retains the low computation cost, introducing only an additional logarithmic term of $k$ and $m$. And when k=mck=m-c ($c$ is a constant), our upper bound strictly matches the lower bound, which is $\Theta\left(c\sqrt{T\log m}\right)$.